OpenStudy (itz_sid):

I NEED HELP PLOX. :3

OpenStudy (itz_sid):

Find the exact length of the curve. #5 Por FAVOR. :3

OpenStudy (itz_sid):

Its my last problem. Yay :D

jimthompson5910 (jim_thompson5910):

Are you able to find the derivative of this?

OpenStudy (itz_sid):

So.. would\[y ^{'}= \frac{ 1-2x }{ 2\sqrt{x-x^2} }+\frac{ \cos^{-1} (\sqrt{x} )}{ 2\sqrt{x} }\]

jimthompson5910 (jim_thompson5910):

that's not a simple sine. That's inverse sine or arcsine

OpenStudy (itz_sid):

Oh right right. Soo... \[y ^{'}=\frac{ 1-2x }{ 2\sqrt{x-x^2} } - \left( \frac{ -1 }{\sqrt{1-x}} \right) \left( \frac{ 1 }{ 2\sqrt{x} } \right)\]

jimthompson5910 (jim_thompson5910):

yes, now let's simplify things \[\Large y ^{'}=\frac{ 1-2x }{ 2\sqrt{x-x^2} } - \left( \frac{ -1 }{\sqrt{1-x}} \right) \left( \frac{ 1 }{ 2\sqrt{x} } \right)\] \[\Large y ^{'}=\frac{ 1-2x }{ 2\sqrt{x-x^2} } +\frac{ 1 }{2\sqrt{x}\sqrt{1-x}} \] \[\Large y ^{'}=\frac{ 1-2x }{ 2\sqrt{x-x^2} } +\frac{ 1 }{2\sqrt{x(1-x)}} \] \[\Large y ^{'}=\frac{ 1-2x }{ 2\sqrt{x-x^2} } +\frac{ 1 }{2\sqrt{x-x^2}} \] \[\Large y ^{'}=\frac{ 1-2x +1}{ 2\sqrt{x-x^2} } \] \[\Large y ^{'}=\frac{ 2-2x}{ 2\sqrt{x-x^2} } \] \[\Large y ^{'}=\frac{ 2(1-x)}{ 2\sqrt{x-x^2} } \] \[\Large y ^{'}=\frac{ \cancel{2}(1-x)}{ \cancel{2}\sqrt{x-x^2} } \] \[\Large y ^{'}=\frac{ 1-x}{ \sqrt{x-x^2} } \] Do you agree with these steps?

OpenStudy (itz_sid):

Oh wait the 1/2ratx wouldnt be needed right?

OpenStudy (itz_sid):

Oh nvm.

OpenStudy (itz_sid):

Yes I agree

jimthompson5910 (jim_thompson5910):

Now plug it into the arc length formula \[\Large L = \int_{a}^{b}\sqrt{1+\left(y^{\prime}\right)^2}dx\]

OpenStudy (itz_sid):

\[\int\limits \sqrt{1+\left( \frac{ 1-x }{ \sqrt{x-x^2} }\right)^2}\]

jimthompson5910 (jim_thompson5910):

yep correct

jimthompson5910 (jim_thompson5910):

the values of 'a' and 'b' aren't given to you, but you can figure them out based on what the domain of f(x) is

OpenStudy (itz_sid):

Um 1?

OpenStudy (itz_sid):

0 and 1?

jimthompson5910 (jim_thompson5910):

correct a = 0 b = 1

OpenStudy (itz_sid):

Oh wait, I did that by looking at the dericative of arctan(sqrt x)

OpenStudy (itz_sid):

How do you do it correctly?

jimthompson5910 (jim_thompson5910):

find the domain?

OpenStudy (itz_sid):

Yes

OpenStudy (itz_sid):

I see the 0, but not the 1

jimthompson5910 (jim_thompson5910):

the first thing that comes to mind is the `sqrt(x)` inside the arcsine the domain for that piece is x >= 0

OpenStudy (itz_sid):

Oh i plugged in a zero for f(x) and got 0.

jimthompson5910 (jim_thompson5910):

The domain of arcsine is -1 <= x <= 1 but we don't have to worry about negatives (since the square root inside won't accept negatives) So far, the domain is 0 <= x <= 1

jimthompson5910 (jim_thompson5910):

The domain of sqrt(x-x^2) is 0 <= x <= 1 as well

jimthompson5910 (jim_thompson5910):

So that's why a = 0 and b = 1

OpenStudy (itz_sid):

When looking for the domain, don't we look to see what numbers make the equation = 0?

jimthompson5910 (jim_thompson5910):

you might be thinking of something like 1/(x-5) there you set x-5 equal to zero and solve to find that x = 5 makes the denominator zero. So we have to kick 5 out of the domain

OpenStudy (itz_sid):

Yea

jimthompson5910 (jim_thompson5910):

we don't have that issue here

OpenStudy (itz_sid):

Oh so we have to kind of visualize the graph?

jimthompson5910 (jim_thompson5910):

yes that's one way to do it. Or to analyze the domain of each piece like I did above

OpenStudy (itz_sid):

Oh okay, I see. So then what would I do after I set up the arc length equation? Do I have to factor the numerator?

OpenStudy (itz_sid):

I get \[\int\limits_{0}^{1} \sqrt{1+\left( \frac{ 1-2x+x^2 }{ x-x^2 } \right)}\]

jimthompson5910 (jim_thompson5910):

\[\Large L = \int_{a}^{b}\sqrt{1+\left(y^{\prime}\right)^2}dx\] \[\Large L = \int_{0}^{1}\sqrt{1+\left(\frac{1-x}{\sqrt{x-x^2}}\right)^2}dx\] \[\Large L = \int_{0}^{1}\sqrt{1+\frac{(1-x)^2}{\left(\sqrt{x-x^2}\right)^2}}dx\] \[\Large L = \int_{0}^{1}\sqrt{1+\frac{1-2x+x^2}{x-x^2}}dx\] Do you see what to do from here?

jimthompson5910 (jim_thompson5910):

ok good, we got the same thing

jimthompson5910 (jim_thompson5910):

now simplify the stuff under the radical further

OpenStudy (itz_sid):

\[\int\limits_{0}^{1} \sqrt{1\left(\frac{ 1-x }{ x } \right)}\]

OpenStudy (itz_sid):

1+*

jimthompson5910 (jim_thompson5910):

Then we can say this \[\Large L = \int_{0}^{1}\sqrt{1+\frac{1-x}{x}}dx\] \[\Large L = \int_{0}^{1}\sqrt{\frac{x}{x}+\frac{1-x}{x}}dx\] \[\Large L = \int_{0}^{1}\sqrt{\frac{x+1-x}{x}}dx\] \[\Large L = \int_{0}^{1}\sqrt{\frac{1}{x}}dx\] which at this point, you can probably see how to evaluate the integral? If not, then here's a hint: \[\Large \sqrt{x} = x^{1/2}\]

OpenStudy (itz_sid):

\[\int\limits_{0}^{1}(x)^{-1/2} = [(x)^{1/2}]_{0}^{1} = 1\]

OpenStudy (itz_sid):

Oops.

OpenStudy (itz_sid):

You also multiply 2. lmao

OpenStudy (itz_sid):

The answer = 2

OpenStudy (itz_sid):

Yay, i got it :D

OpenStudy (itz_sid):

Thanks!

jimthompson5910 (jim_thompson5910):

yep the final answer is 2